Ten competitors in a beauty contest are ranked by three judges in the following order:-

First Judge

1 6 5 10 3 2 4 9 7 8

Second Judge

3 5 8 4 7 10 2 1 6 9

Third Judge

6 4 9 8 1 2 3 10 5 7

a. Use the rank correlation co efficient to discuss which pair of judges have the nearest approach to common tastes in beauty.

b. Could this problem be studied using correlation if yes then how? If not then why? Explain.
1
Expert's answer
2020-08-11T19:14:07-0400

Here we calcuate three rank correlation coefficient:

r12 beteen the ranks of judges I and II

r23 beteen the ranks of judges II and III

r13 beteen the ranks of judges I and III



r12=1-(6"\sum"D212)/(N(N2-1))=1-(6*200)/(10*99)=-0.212

r23=1-(6"\sum"D223)/(N(N2-1))=1-(6*214)/(10*99)=-0.297

r13=1-(6"\sum"D213)/(N(N2-1))=1-(6*60)/(10*99)=0.636

The Judges I and III have nearest approach to common tastes in beauty where as the Judges II and III disagree the most

b. yes, to find out the nearest approach to common tastes in beauty judges, we should calculate the sets of correlation between the pairs of judges